OpenStudy (anonymous):

find the rate of change of the line containing (5,-2) and (5,6)

OpenStudy (hba):

\[m=\Delta y /\Delta x = y_2-y_1/x_2-x_1\]

OpenStudy (hba):

My interpration maybe wrong :/

OpenStudy (anonymous):

simply find the slope thats the rate of change

OpenStudy (hba):

That is the formula for the slope.

OpenStudy (hba):

@Judith5 Can you solve it ?

OpenStudy (hba):

(x1,y1)=(5,-2) (x2,y2)=(5,6)

OpenStudy (anonymous):

would i divide next or would (5,-2) and (5,6) be my answer?

OpenStudy (hba):

Well i gave you a formula m=y2-y1/x2-x1 plug in the values and find your answer

OpenStudy (anonymous):

oh ok thank you(:

OpenStudy (hba):

Your'e welcome :)

OpenStudy (agent0smith):

The x-value hasn't changed, the slope is undefined. At x=5, y = -2, and also at x=5, y=6 Trying to find the gradient will end like this: m = (6-(-2))/(5-5) = 8/0

OpenStudy (agent0smith):

 The line is a vertical line, x=5.